### v1

# An FPGA-Based Open-Source Hardware-Software Framework for Side-Channel Security Research

Abstract
Attacks based on side-channel analysis (SCA) pose a severe security threat to modern computing platforms, further exacerbated on IoT devices by their pervasiveness and handling of private and critical data. Designing SCA-resistant computing platforms requires a significant additional effort in the early stages of the IoT devices’ life cycle, which is severely constrained by strict time-to-market deadlines and tight budgets. This manuscript introduces a hardware-software framework meant for SCA research on FPGA targets. It delivers an IoT-class system-on-chip (SoC) that includes a RISC-V CPU, provides observability and controllability through an ad-hoc debug infrastructure to facilitate SCA attacks and evaluate the platform’s security, and streamlines the deployment of SCA countermeasures through dedicated hardware and software features such as a DFS actuator and FreeRTOS support. The open-source release of the framework includes the SoC, the scripts to configure the computing platform, compile a target application, and assess the SCA security, as well as a suite of state-of-the-art SCA attacks and countermeasures. The goal is to foster its adoption and novel developments in the field, empowering designers and researchers to focus on studying SCA countermeasures and attacks while relying on a sound and stable hardware-software platform as the foundation for their research.
Index Terms:
Side-channel analysis, RISC-V, open source, open hardware, field-programmable gate array, system-on-chip, Internet of things, research.
I Introduction
Hundreds of billions of Internet of Things (IoT) devices are getting more and more able to autonomously make decisions thanks to artificial intelligence (AI). As they get more capable, pervasive, connected, and able to interact with each other through 5G and 6G networks, IoT devices are constantly collecting, processing, and exchanging personal, private, and critical data, making it paramount to consider security as an essential aspect during their design and whole life cycle [1].
Relying on the theoretical security provided by traditional cryptography solutions such as AES, RSA, and SHA-3 and by secure communication protocols such as TLS and SSH is, however, insufficient to guarantee IoT device security. Side-channel analysis (SCA) attacks, in particular, target information collected from executing a specific implementation of a cryptographic scheme or security protocol rather than flaws in their specification or the theoretical problems at their foundation [2]. Such so-called side-channel information ranges from the time required for a cryptography-related computation to the power consumption, electromagnetic emissions, and noise produced during its execution. Moreover, while some fault-injection-based attacks interfere with the correct functioning of the target device, many other so-called passive side-channel attacks, such as differential power analysis (DPA) ones, are undetectable by the system under attack.
While SCA attacks increasingly become serious security threats to IoT devices, and in particular to those that operate in public spaces and are accessible by anyone, the computing platforms that power these devices and that handle personal, private, and critical data are, however, often not designed to protect from them. The commercially available microcontroller-based system-on-chips (SoCs) that provide the computing and connection capabilities of the IoT devices indeed commonly execute cryptography applications, while not being meant to do so in a secure way, mostly due to cost-saving reasons [3].
At the same time, the strict time-to-market deadlines, tight energy and power consumption and area constraints, and rapidly evolving tasks to be executed on such SoCs make it harder to offload their tasks, including cryptographic ones, to dedicated hardware accelerators and favor instead the design of simpler and more flexible programmable architectures that can also run an operating system (OS).
Conversely, the open literature does not provide a comprehensive hardware-software solution to tackle the challenge of SCA security on modern RISC-V platforms. The few RISC-V security-oriented solutions, e.g., OpenTitan [4], are not meant as general-purpose computing platforms and do not support the development of novel SCA attack and defense methodologies, while others focus their research effort solely on the hardware side while disregarding the software framework [5] or viceversa.
As new SCA attacks continue to emerge, designing novel IoT devices that are secure against them requires providing the system designers and security researchers with a sound and stable hardware-software platform that lets them focus on analyzing the security against SCA attacks and developing countermeasures to thwart the latter. Moreover, there is a need to not only detect the presence of vulnerabilities, i.e., of so-called information leakage, but also to pinpoint the specific signals in the hardware design that cause such leakage, enabling the designers with the possibility to more easily correct their systems and make them resistant to SCA attacks.
Contributions
This manuscript introduces an open-source research framework for side-channel analysis on FPGA-based IoT/microcontroller-class computing platforms, that encompasses a complete SoC and a software toolchain for SCA attacks and countermeasures, delivering a comprehensive hardware-software solution that is, to the best of our knowledge, currently lacking in the open literature.
The framework provides an IoT-class SoC that includes a CPU based on open-standard RISC-V as its instruction set architecture (ISA), dedicated hardware to enable the implementation of state-of-the-art SCA countermeasures, an ad-hoc debug infrastructure to maximize the observability and controllability of the computing platform and thus simplify the execution of SCA attacks, and support for the open source FreeRTOS real-time operating system (RTOS). A complete flow encompasses the configuration of the SoC, the execution of target applications and corresponding collection of side-channel information, and the analysis to identify SCA vulnerabilities and leakage sources.
The goal of empowering designers and researchers to only focusing on studying SCA countermeasures is achieved through three main contributions:
  • Capability to identify leakage sources – Hardware security requires not just identifying the presence of a vulnerability, but also pinpointing the source of such leakage. The computing platform has a minimal architecture to expose the least leakage sources and make them eventually simpler to identify, in addition to being easier to emulate and cheaper in terms of area and power cost, while a one-to-one match between the prototyped platform and its emulated counterpart is enforced through dedicated hardware mechanisms to enable the collection of the most accurate side-channel signal.
  • Usability for research purposes – The framework is released open source, including all the hardware and software, from the SoC to the software scripts driving the SoC configuration, compilation, prototype execution and emulation, oscilloscope measurement, and state-of-the-art SCA attacks, to foster its adoption in research settings. Carrying out an analysis through the proposed framework provides not only the traditional power, performance, and area (PPA) quality metrics, but it extends them by adding the security dimension.
  • Complete SCA ecosystem – The framework includes dedicated hardware and software support for state-of-the-art SCA attacks and countermeasures that can be employed out of the box to directly compare with solutions from the literature as well as evaluate and enforce the SCA security of the computing platform. The experimental evaluation in this manuscript demonstrates the framework’s capabilities and showcases the SCA attacks and countermeasures included as part of its open-source release.
Structure of the manuscript
The rest of this manuscript is organized into five parts. Section II provides a background on SCA attacks and countermeasures and an overview of the state-of-the-art RISC-V computing platforms. Section III describes the proposed framework and the hardware-software infrastructure that implements it. Section IV details the key hardware and software aspects that enable the SCA attack and countermeasure capabilities of the framework, Section V showcases the framework capabilities through a comprehensive experimental evaluation. Finally, Section VI draws conclusions and discusses the future works and developments on the proposed framework.
II Background and Related Works
II-A Side-channel analysis attacks and countermeasures
A vast amount of research has tackled the complementary topics of SCA attacks and countermeasures ever since the emergence of SCA as a security threat [6].
SCA attacks can be classified as either non-profiled or profiled. Non-profiled ones, such as those based on differential power analysis (DPA) [6] and correlation power analysis (CPA) [7], only target side-channel information that is obtained from the specific device under attack while it is performing a computation, and then attempt to recover the corresponding secret key by leveraging the statistical correlation between the measured side-channel signal and the data being processed, exploiting a partial knowledge of the latter. Profiled attacks [8] are carried out instead by initially making use of a replica of the device to be attacked to identify and fine-tune the side-channel leakage model and later employing such model to attack the actual target device, under the assumption that the latter and its replica share identical or at least similar leakage models [9]. Whereas SCA attacks have traditionally employed statistical techniques, machine- and deep-learning approaches recently emerged as promising research avenues for more capable attacks, ranging from the identification of cryptographic operations from side-channel power traces [10] to learning-based non-profiled [11] and profiled [12] attacks.
The advancements in SCA attacks are mitigated by an even larger research effort being devoted to identifying new defense mechanisms against them. SCA countermeasures can be mainly split into masking and hiding ones. Masking countermeasures split the sensitive intermediate values into different shares that are computed independently from each other, with the goal of minimizing the dependency of each share from the secret key [13]. Hiding countermeasures aim instead to randomize or add noise to the side-channel emission of the computing platform to be protected in order to reduce the information leakage that can be exploited by an attacker. Such countermeasures make use of techniques such as code morphing, i.e., the insertion of random instructions in the original execution flow [14], clock frequency randomization [15], and the concurrent computation of multiple cryptographic operations with invalid keys [16].
Figure 1: High-level flow of the proposed framework.
II-B State-of-the-art RISC-V computing platforms
RISC-V has emerged as the de-facto standard ISA for IoT-class computing platforms thanks to its modularity, enabled by a minimal baseline instruction set whereas separate ISA extensions are each devoted to a specific set of additional features, and extensibility, given by the possibility of exploiting unused opcode space to encode custom instructions and extensions according to the specific needs of computer designers [17].
On the one hand, a variety of RISC-V-based microcontroller-class SoCs is available, such as PULPino [18] and GAP-8 [19], while ultra-low-power platforms also leveraged RISC-V in multi-core heterogeneous architectures, e.g., Mr. Wolf [20] and HERO [21]. Such solutions focus on delivering good performance at a low energy and power consumption, but do not cover security and in particular SCA-related aspects.
On the other hand, frameworks such as ESP [22] and Chipyard [23] have emerged in the literature for the agile development of heterogeneous multi- and many-core SoCs that feature both general-purpose CPU cores and hardware accelerators [24]. However, they focus on the modularity and composability of the obtained computing platforms and on the fast prototyping pf the latter on FPGA targets, while they do not consider instead any security aspects.
All the previously listed platforms, including microcontroller-class, ultra-low-power, and heterogeneous multi- and many-core SoCs, are capable and efficient on the computing side. However, their complexity hinders the possibility of simulating them in an RTL simulator, which is crucial to detecting SCA vulnerabilities and, more importantly, accurately identifying which signals are responsible for the information leakage, which becomes ever more challenging to achieve as such hardware platforms keep growing in size.
Finally, the OpenTitan project [4] delivers a hardware root of trust (RoT) centered around a RISC-V core, not providing instead neither a complete SoC platform nor a hardware-software framework for SCA-related research.
There also exists a variety of both RISC-V resource-constrained and application-class cores in the open literature, including CORE-V CV32E40P [25] and CVA6 [26], UCB’s Rocket [27], and Frontgrade Gaisler NOEL-V [28], that are however of limited use, when considered on their own, in providing a full platform for SCA research. State-of-the-art architectural simulators such as gem5 [29] and RISC-V-specific GVSoC [30] operate instead at a high level of abstraction, drastically shrinking the time required to simulate a computing platform but conversely making them unsuitable to produce information that can be effectively exploited by SCA attacks.
Conversely, a large part of the literature in the SCA research field targets instead more outdated hardware such as the STM32 32-bit Arm Cortex-M [31] and ATmega 8-bit AVR [32] microcontrollers, highlighting the need to foster novel SCA-related research on modern IoT-class computing platforms.
III Framework
Figure 2: Hardware-software infrastructure that implements the proposed framework.
The proposed framework enables thoroughly evaluating the vulnerability to SCA attacks of cryptographic applications executed on an IoT-class SoC that implements countermeasures through dedicated hardware support. The possibility to accurately match the emulated and prototype execution of the computing platform allows not only detecting information leakage, but even more importantly to identify its sources, which is paramount to fixing the vulnerabilities of the SoC during the design phase. Section III-A outlines the high-level flow for the proposed framework, while Section III-B discusses how such flow is implemented in a hardware-software infrastructure.
III-A High-level flow
The proposed framework, depicted in Figure 1, can be described as a sequence of three main phases devoted to 1) configuringthe hardware-software setup, 2) measuringSCA-related information from simulation and prototype execution, and 3) analyzingsuch information to detect leakage and identify its sources. Remarkably, the open-source nature of the framework, as well as the adoption of standard languages for all the inputs and its internals, enable the user to tinker, adapt, and tailor it to their needs. Software scripts automate the execution of each of the three phases, which are discussed in more detail in the rest of this section.
III-A1 Configure
The high-level flow starts with the configuration of the hardware and software parts of the target to be simulated and prototyped. The SoC is thus configured by the user and implemented according to the standard hardware design flow, while the applications and, optionally, the RTOS are compiled to produce an executable to be run by the CPU.
Implement
The framework provides the RTL description of the SoC and the design constraints files that define the frequency of the clock signal and map the I/O to the target FPGA chip, thus enabling the deployment on the prototype board. The initial step requires selecting which additional components to instantiate in the SoC, the parameterization for configurable aspects of the SoC, and which FPGA to target, then an EDA toolchain is leveraged for the synthesis and place-and-route of the netlist and the generation of the corresponding FPGA bitstream.
Compile
During this phase, the source code for the applications provided by the user of the proposed framework is compiled, through a compiler toolchain for RISC-V, to be executed on the target platform, both in simulation and on the prototype. The compilation produces an executable file for the compiled applications and, optionally, the RTOS, meant to be executed by the computing platform both in simulation and on the prototype FPGA. The boundaries for the time window of interest in the execution of the application, e.g., the computation of a specific cryptographic kernel of which to evaluate SCA resistance, are also extracted from the executable file.
III-A2 Measure
Once the netlist and bitstream for SoC have been generated and the applications have been compiled, the simulation and prototype execution can be carried out to collect the measurements meant to be used for SCA purposes. The simulation produces switching activity statistics, while a power trace is measured from an oscilloscope connected to the board during the prototype execution. The simulation and prototype execution are notably time-aligned, i.e., synchronized, through dedicated hardware features implemented in the target SoC platform. Such synchronization enforces indeed a temporal match between the switching activity obtained from simulation and the power trace measured during the prototype execution.
Simulate
The simulation, in a SystemVerilog testbench provided as part of the framework, of the post-place-and-route netlist of the SoC running the compiled application executable produces the switching activity of the internal signals of the SoC in the time window of interest.
Execute
After flashing the bitstream to the FPGA mounted on the prototype board, the SoC is fed the application binary to be loaded into memory and then executes it. An oscilloscope collects the power trace measurement for the execution of the target application on the prototype board within the time window of interest, matching the one employed in the corresponding simulation.
III-A3 Analyze
The final phase of the proposed framework foresees the analysis of the measurements from the previous phase, i.e., the switching activity of the simulation and the power consumption trace of the board execution, both corresponding to the time window of interest, in order to detect whether there was any cryptographic information leakage as well as to identify the eventual sources of such leakage. A set of state-of-the-art SCA techniques, ranging from CPA and template to ML-based ones, is therefore provided to evaluate the SCA security of the platform and the application executed on top of it. Remarkably, the analysis phase also produces the traditional power, performance, and area (PPA) metrics.
III-B Hardware-software infrastructure
The hardware-software infrastructure depicted in Figure 2 realizes the high-level flow described in Section III-A by means of three main components, namely 1) a (virtual) prototype, 2) an oscilloscope, and 3) a host PC. This part discusses in detail the three components and how they interact with each other to deliver the proposed framework.
III-B1 (Virtual) prototype
The computing platform at the foundation of the proposed SCA research framework implements an SoC architecture that can notably be instantiated as a prototype on FPGA, to collect the power trace measurements from an oscilloscope, as well as emulated as a virtual prototype in a RTL simulator that instantiates its post-place-and-route netlist in a testbench, to generate the matching VCD switching activity. The correspondence between the prototype and its virtual counterpart, in particular in how they execute a target application, is enforced by dedicated hardware mechanisms and it is the crucial aspect that enables employing our framework for SCA analysis.
The SoC architecture is built around a Wishbone interconnect. It comprises a CPU, a dynamic frequency scaling (DFS) actuator, and a global debug unit as its masters and a memory, a user UART, a true random number generator (TRNG), and a timer as its bus slaves.
The single-core, in-order, five-stage pipelined, 32-bit RISC-V CPU implements the base integer (I) and integer multiplication and division (M) RISC-V 32-bit ISA extensions [33]. Supporting only I and M extensions, i.e., the bare minimum extensions for an IoT-class CPU, provides the simplest and most observable setup, making it easier to carry out the side-channel analysis. A floating-point unit (FPU) can however be optionally instantiated as a functional unit of the CPU to expand the computing capabilities of the SoC [34, 35]. The interconnect consists of two separate 64-bit Wishbone data and instruction buses, both supporting single read and write transactions as well as burst ones, according to a modified Harvard architecture to minimize contention. The DFS actuator enables changing at run time the frequency of the clock signal fed to the CPU, whereas the rest of the SoC operates at a fixed clock frequency. A global debug unit and local debug units connected to the former in a point-to-point fashion compose the debug infrastructure of the SoC. The global debug unit interacts with the host PC through the system UART while two different channels support its communication with the rest of the SoC. Memory location of the memory-mapped bus slaves are accessed through Wishbone reads and writes, while dedicated lines connect the global debug unit to the two local ones that manage the CPU and DFS actuator, respectively.
The main memory, making use of the block RAM (BRAM) resources available on AMD FPGAs, is instantiated as a slave on the Wishbone data bus. The bus slaves include a TRNG, that produces a sequence of random bits meant to be used in cryptography and SCA countermeasure tasks, a timer, which enables support for FreeRTOS, and a user UART, that provides an I/O interface for the application. Hardware accelerators exposing a Wishbone interface can also notably be added as memory-mapped bus slaves to the SoC to extend its capabilities.
The implementation phase of the proposed flow includes the selection of which parts of the SoC to instantiate, e.g., the optional DFS actuator, TRNG, and timer, and the configuration of its parametric features, e.g., the baud rate of the UART interface, the width of the instruction and data buses, and the branch prediction scheme of the CPU.
In addition to the UART I/O interfaces, the prototype exposes the FPGA voltage (VDD in Figure 2), which is related to the power consumption of the whole FPGA chip, and a trigger signal (Trg), that is driven by the local debug unit of the CPU.
III-B2 Oscilloscope
Power measurements from the prototype execution are carried out through an oscilloscope with two analog channels and a frequency bandwidth that is sufficient to collect samples from the target platform under measurement without aliasing. The oscilloscope is connected to the prototype board, with an analog channel measuring the voltage of the FPGA and the other one monitoring a signal meant to trigger the data acquisition on the former. It is managed through a USB interface by the host PC, which takes care of its configuration and receives the data samples measured from the board on both analog channels.
Figure 3: Detailed microarchitecture of the debug subsystem, timer, and TRNG.
III-B3 Host PC
The host PC drives the whole proposed framework through the flow software scripts, which manage the interaction with both hardware devices, namely the prototype board and the oscilloscope, and software tools, such as the EDA and compilation toolchains and state-of-the-art SCA attack scripts. Matching the high-level flow previously described in Section III-A, we can identify three main top-level scripts that drive the corresponding phases of the framework.
Configure
The configuration script leverages an EDA toolchain for the synthesis and place-and-route and bitstream generation of the SoC, previously configured according to the needs of the user, also taking as its inputs the XDC constraint files that include information about the internal clock signals as well the I/O mapping on the target FPGA chip. A RISC-V compiler toolchain is employed instead to compile the applications and, optionally, the RTOS. The compilation of the sources can thus include the sources for FreeRTOS, when targeting the execution of applications on top of the RTOS, or not, when the goal is instead bare-metal execution. A linker script and startup routines to be executed before the program’s main function are included to support the latter case. The compilation process produces an executable file, that can be loaded into the memory of the computing platform to be executed both in its simulation and on the prototype FPGA. Boundaries for a time window of interest are obtained from the executable to enable accurately matching, also from a temporal standpoint, the simulation of the computing platform with its prototype execution.
Measure
The simulation leverages an RTL simulator to simulate the post-place-and-route netlist of the SoC inside a SystemVerilog testbench provided as part of the framework. The testbench loads the VMEM for the target application into the memory of the target SoC and drives the execution of the application according to the time window boundaries extracted from its ELF file at the previous phase. The simulation outputs a value change dump (VCD) that contains all the switching activity of the internal signals of the SoC corresponding to the execution of the application within the time window of interest The prototype execution requires, after flashing the bitstream to the FPGA mounted on the prototype board, feeding the application binary to the SoC to be loaded into its memory. The flow script properly drives the board execution through the SoC debug interface, matching the behavior of the SystemVerilog testbench, and manages the oscilloscope connected to the board to collect the power trace of the execution. The VCD from the simulation and the power trace from the prototype execution are saved to a data storage device, so that they can be later retrieved for performing the various analyses.
Analyze
The VCD obtained from the simulation and the power traces collected from the prototype execution, both corresponding to the exact same time window, are retrieved from the storage where they were saved to analyze them through a set of state-of-the-art SCA techniques, including CPA, template, and ML-based attacks, to detect whether there was any cryptographic information leakage and also identify the eventual sources of the latter. In addition to the SCA security statistics, the analysis phase outputs the traditional PPA metrics leveraging the power consumption measured from the board execution, the cycle-granularity latency collected from simulation, and the resource utilization reports from netlist implementation. The outputs of the computation can also be obtained through the debug infrastructure to further check the correct functioning of the system.
IV Microarchitecture
Dedicated hardware and software mechanisms enable the SCA attack and countermeasure capabilities of the proposed framework. The debug subsystem, with its breakpointing and ad-hoc triggering mechanism, ensures maximum observability and controllability of the computing platform, thus allowing the collection of accurate side-channel information that can then be exploited by SCA attacks of different kinds. A TRNG, a DFS actuator, and a timer can instead be optionally instantiated in the SoC to enable support for a variety of state-of-the-art SCA countermeasures ranging from masking to hiding ones. The rest of this section delivers a detailed discussion of the microarchitecture of the debug subsystem, TRNG, DFS actuator, and timer.
IV-A Debug subsystem
The framework provides maximum observability and controllability of the computing platform through the hardware support for breakpoints and triggerpoints as well as a dedicated debug infrastructure that exposes a GDB-like debug interface, thus enabling the collection of accurate side-channel information that can then be targeted by SCA attacks of different kinds. This part discusses first the global-local debug microarchitecture, then focuses on the breakpointing and triggering mechanisms, and finally provides an overview of the debug capabilities that can be exploited by the external host PC.
IV-A1 Global and local debug units
The debug subsystem enables controlling and observing the whole SoC through a message-based protocol. It is composed of a global debug unit and of a number of local debug units. The global one, that acts as a master on the Wishbone bus, receives debug messages through a system UART interface and accordingly communicates both with the other bus masters, through point-to-point connections to local debug units that act as adapters to interface with such masters, and with the bus slaves, with whom it interacts directly through the bus by exploiting their memory-mapped nature.
Two local debug units are instantiated for the CPU core and for the DFS actuator, respectively, to act as adapters between the global debug unit and the two bus masters. The local debug unit interface with the global one is common to all the local debug units, while the interface with the CPU or DFS actuator is custom tailored to the specific interactions that are implemented with such module. The CPU local debug unit can access the program counter (PC), registers, and performance monitoring counters (PMCs) of the CPU, as well as halting, resetting, and restarting it and advancing its execution by a single step. The DFS one can set a new target clock frequency for the DFS actuator, get the frequency of the current clock signal generated and configure the DFS to randomly switch the frequency of the clock signal. Figure 3 depicts the microarchitecture of the debug subsystem of the SoC, highlighting the global debug unit, the CPU and DFS local debug units, and their interactions with each other as well as with the CPU and the DFS actuator.
(a) Debug messages
(b) Command field
TOKEN_TYPE R/W Destination Group
INVALID Read Memory- Invalid
mapped
MMAP_READ Read Memory- Memory-
MMAP_WRITE Write mapped mapped
GPR_INT32_READ Read CPU local CPU
GPR_INT32_WRITE Write debug unit registers
GPR_FPU32_READ Read
GPR_FPU32_WRITE Write
HALT_CPU Read CPU local CPU reset/
RUN_CPU Read debug unit resume/halt
RST_CPU Read
GET_DULOCAL_STATE Read
GET_CPU_PC Read
ADVANCE_ONE_STEP Read CPU local CPU
ECHO_FRONTEND Write debug unit stepping
GET_LOW_CYCLECNT Read CPU local Performance
GET_HIGH_CYCLECNT Read debug unit counters
GET_LOW_INSTRCNT Read
GET_HIGH_INSTRCNT Read
SET_BRKPNT_CPU Write CPU local Breakpoints
GET_BRKPNT_CPU Read debug unit configuration
RM_BRKPNT_CPU Write
GET_NUM_BRKPNT_CPU Read
SET_TRGPNT_CPU Write CPU local Triggerpoints
GET_TRGPNT_CPU Read debug unit configuration
RM_TRGPNT_CPU Write
GET_NUM_TRGPNT_CPU Read
SET_FREQ_DFS Write DFS local DFS
GET_FREQ_DFS Read debug unit configuration
RND_FREQ_DFS Write
(c) Token types
Figure 4: Debug messages supported by the proposed framework: (a) structure and width of the debug messages, (b) encoding of the command field, (c) token types of the request messages. Legend: BTE burst type extension, W write enable (0: read, 1: write), CTI cycle type identifier (for burst mode), SEL select, TOKEN_TYPE request message type, Reserved reserved for future use, CPU_ID identifier for target local debug unit, A ack, E error; BTE, W, CTI, and SEL refer to Wishbone.
IV-A2 Breakpoints and triggerpoints
The CPU local debug unit supports, through its interaction with the CPU, a traditional breakpointing system coupled with an ad-hoc triggering one. Breakpoints enable halting the CPU when the PC matches their corresponding addresses, while triggerpoints, an ad-hoc variant of breakpoints, are meant to toggle a trigger signal that drives the data acquisition from the oscilloscope and do not instead halt the CPU, which is kept regularly running. The local debug unit for the CPU contains two tables for the breakpoint and triggerpoint addresses, respectively, as shown in Figure 3. Both tables include a configurable number of entries, each composed of a 32-bit address that corresponds to an instruction in the target application and a 1-bit flag that signals its validity.
The valid addresses in the breakpoint table are constantly checked against the current PC of the CPU, whose value is passed to the CPU local debug unit. Whenever there is a match between a valid breakpoint address and the current PC, the CPU is halted until it is resumed through a dedicate debug command. The triggerpoint table entries are similarly compared against the PC, but a match between the latter and a triggerpoint address produces a notably different effect. Rather than halting the CPU as in a traditional breakpoint fashion, reaching a triggerpoint toggles a 1-bit signal that is mapped on an I/O pin of the prototype board to be used as a trigger signal for an oscilloscope.
IV-A3 Debug messages
The debug infrastructure exposes a request-response protocol to interface with the SoC through the system UART. The communication is indeed carried out as a sequence of request and response messages. Notably, no new request message can be issued until the previous request has been completed and the corresponding response message has been sent back. Figure 4 summarizes the width, structure, and information encoded in the various messages depending on whether they are request or response and read or write and on whether they are intended for memory-mapped recipients or local debug units.
Request messages can be of four types, depending on whether they correspond to read or write actions and whether their recipients are memory-mapped devices or bus masters. All request messages are composed of a 16-bit command and a 32-bit, while write requests also comprise a 32-bit data, as shown in Figure 4(a). The command field encodes different information depending on whether the debug message targets a memory-mapped peripheral or a local debug unit, as depicted in Figure 4(b). Debug request messages that can be sent to the SoC through the system UART include, as listed in Figure 4(c), i) memory-mapped reads and writes, ii) register reads and writes, iii) commands to reset, resume, and halt the CPU, advance its execution by a single step, and get its current state and program counter, iv) commands to retrieve performance monitoring counters v) commands to set, get, and remove both the breakpoints and the triggerpoints, and vi) commands to set and get the target clock frequency for the DFS actuator and configure it to randomly switch clock frequencies.
Response messages can be instead of only two types, i.e., either read or write ones. The former include data as part of the response, while the latter only acknowledge the completion of the requested operation or the occurrence of an error.
Figure 5: Detailed microarchitecture of the DFS actuator.
IV-B TRNG
The TRNG is a Wishbone slave peripheral that exposes a set of 32 random bits through a memory-mapped register. The TRNG can be configured in the architecture of the digital noise sources and post-processing methods that compose it to obtain different results in terms of FPGA resource utilization, throughput, and security. A digital noise source produces the actual entropy underlying the random number generation, while a post-processing method improves the statistical and security properties of the TRNG. Three digital noise sources and three post-processing methods from the literature are provided with the proposed framework to implement the TRNG component. The digital noise sources that can be instantiated in the TRNG module are NLFIRO, PLL-based, and edge-sampling ones, and they can be coupled with XOR, Von Neumann, and LFSR post-processing methods [36]. The random output produced by the TRNG component is periodically refreshed and exposed by a memory-mapped register that can be read through Wishbone.
IV-C DFS actuator
The DFS actuator, whose microarchitecture is depicted in Figure 5, leverages two mixed-mode clock manager (MMCM) components to provide a glitch- and latency-free switching between different clock frequencies at run time. It receives as its inputs a target clock frequency (fin in Figure 5), a 1-bit flag to enable random DFS (rnd), and a reference clock signal (clkref), and it outputs the generated clock signal (clkout) and the current clock frequency (fout). The rnd flag selects the actual target clock frequency for the DFS reconfiguration between the one received through the fin input by the DFS local debug unit and a random value frnd output by an internal TRNG with similar to the one in Section IV-B. Such target clock frequency, registered to be output to the DFS local debug unit and readable through a dedicated debug command, is decoded and used as an address to select a corresponding set of MMCM configuration parameters from the BRAM memory.
Notably, the clock signal output by the MMCMs of an AMD FPGA remains low during its reconfiguration, thus causing a clock-gating effect on the computing platform. The proposed DFS actuator avoids such negative effect by employing two MMCM components (MMCMA and MMCMB in Figure 5), among which the FSM logic selects a master and a slave. The master MMCM keeps generating the output clock signal while the slave one is under reconfiguration, after which their roles are swapped. An ack signal monitors the locking status (locki) of the two MMCMs to prevent triggering a new reconfiguration through the cfgi flag while the previous one is still in progress. The clkout clock output by the DFS actuator is selected through the isMst 1-bit flag as the one output by the current master between the clkoutA and clkoutB clock signals output by the two MMCMs, respectively.
The flexible design of the DFS architecture supports up to 1024 different clock frequency configurations. The BRAM memory can indeed store up to 1024 user-defined sets of MMCM configuration parameters, with a minimum step of 0.1250.1250.1250.125MHz and lowest and highest achievable clock frequencies of 5555MHz and 800800800800MHz, respectively, due to MMCM feasibility constraints. Random reconfiguration is activated through a debug request (RND_FREQ_DFS in Figure 4(c)), which is dispatched to the DFS local debug unit that in turn asserts the rnd input flag. When rnd is set to 1, the MMCMs keep indeed reconfiguring to new clock frequencies as soon as the previous reconfiguration has been completed, i.e., the ack signal to the FSM is asserted.
IV-D Timer and FreeRTOS support
The timer is a memory-mapped peripheral that exposes two 64-bit registers which can be accessed through read and write requests on the Wishbone interconnect. The two registers correspond to the current time, which gets increased by 1 at each clock cycle, and to the time threshold. Once the current time is greater than the time threshold, an interrupt request to the CPU is raised by setting to 1 the content of a 1-bit register.
Instantiating a timer in the SoC notably enables executing the FreeRTOS real-time operating system and thus providing support for implementing coarse-grained multithreading.
V Experimental Evaluation
The framework is released open source, including all the hardware and software that comprise it. The open-source nature of the framework, as well as the adoption of standard languages ranging from SystemVerilog for the RTL description of the SoC and the testbench to C for the application and RTOS sources and Python for the flow scripts, enable the user to expand its capabilities and support novel SCA attacks and countermeasures. The goal of the proposed framework is not introducing new SCA attack or countermeasure techniques, but providing instead an infrastructure that implements the current state of the art and thus enables new research in the field. The experimental campaign described in this section aims to showcase how the proposed framework delivers an effective research platform for SCA attacks and countermeasures on FPGA targets. Section V-A lists the hardware and software required by the framework, Section V-B details the specific setup employed for the experimental evaluation, Section V-C delivers an experimental analysis of the SCA techniques provided out of the box.
V-A Hardware and software requirements
The framework makes use of widely available devices and tools, including the development board and FPGA chips for prototyping, the digital oscilloscope for power trace measurement, and the software required for the various phases of the flow ranging from EDA synthesis and place-and-route to compilation and SCA resistance evaluation.
V-A1 Prototype
XDC design constraint files are provided for prototyping on the NewAE Technology CW305 Artix FPGA Target board, specifically designed for power analysis and fault injection attacks against hardware cryptographic functions [37]. The CW305 board mounts a BGA socket that can accommodate any chip from the AMD Artix-7 mid-range FPGA family, which is the most widely adopted in academic and industrial research, including the security and cryptography fields [38]. The proposed SoC fits into AMD Artix-7 50 chips, which pack 32600 look-up tables (LUT), 65200 flip-flops (FF), 120 digital signal processing (DSP) blocks, and 75 36kb blocks of block RAM (BRAM), and larger ones from the same family. Simulation of the virtual prototype is supported by the AMD Vivado EDA toolchain.
V-A2 Oscilloscope
Power measurements are carried out through an oscilloscope from the Pico Technology PicoScope 5000 Series family, which features a 200MHz maximum bandwidth and can collect up to 1 billion samples per second at a resolution ranging from 8 to 16 bits [39]. The oscilloscope has its A analog channel connected to the board through a coaxial SMA connector to measure the FPGA voltage, while its B analog channel is connected to a general-purpose I/O pin that outputs the trigger signal which marks the start and the end of the computation.
V-A3 Host PC
The host PC must be able to execute the AMD Vivado toolchain for the synthesis, place-and-route, and bitstream generation targeting AMD Artix-7 FPGAs, as well as for the simulation and switching activity VCD collection. The host PC is also in charge of the compilation of the application sources, optionally together with the FreeRTOS ones, through a C compiler toolchain for RISC-V targets. Drivers and software for the CW305 board and PicoScope 5000 oscilloscope, connected to the host PC through USB interfaces, are required to drive the prototype execution and power trace collection. All the VCD, power trace, area, and execution time data is stored permanently on a disk mounted on the host PC to be used in the analysis phase. Support for Python 3.9 or higher is necessary to run the software scripts that manage the whole framework as well as the scripts that implement state-of-the-art SCA attack techniques.
V-B Experimental setup
V-B1 Software setup
The framework is run in Ubuntu 22.04.3 LTS on a host PC that features an Intel i7-10700 CPU and a 64GB DDR4 memory. The host PC includes ChipWhisperer 5.1.0 software and PicoScope 7 software and drivers to manage the CW305 board and the oscilloscope, respectively. Applications are compiled with GCC 11.4.0 to be executed on the SoC either bare-metal or on top of FreeRTOS 11.0.1. AMD Vivado ML 2023.1 is employed for the RTL synthesis, place-and-route, bitstream generation, and simulation.
TABLE I: Breakdown of the SoC’s FPGA resource utilization.
FPGA resource utilization
Component LUT $F$ DSP BRAM
CPU 4386 3192 4 0
DFS 822 225 0 1.5
Global debug unit 286 248 0 0
CPU local debug unit 875 587 0 0
DFS local debug unit 19 29 0 0
Memory 206 138 0 64
TRNG 2486 1252 0 0.5
Timer 144 163 0 0
System UART 365 292 0 0
User UART 359 284 0 0
Overall SoC 10667 7563 4 66
V-B2 Hardware setup
The FPGA target in the experimental evaluation is an Artix-7 100 (xc7a100tftg256-1) chip mounted on a CW305 board. The oscilloscope is a PicoScope 5244D, that features two analog channels and a 200MHz bandwidth. The SoC to be prototyped on FPGA and simulated is configured with a CPU that implements the RV32IM instruction set, a 256kB main memory, and a TRNG, a DFS actuator, and a timer that enable support for all the included state-of-the-art SCA countermeasures. TABLE I lists the FPGA resource utilization resulting from synthesis and place-and-route targeting 100 MHz and 50MHz clock frequencies for the CPU, driven by the DFS actuator, and the rest of the SoC.
V-B3 Breakpoints and triggerpoints
Two triggerpoint addresses are assigned, respectively, the addresses of instructions in the target application between which to acquire the power measurement from the oscilloscope. In particular, reaching the start triggerpoint address raises a trigger signal mapped on an I/O pin of the prototype board, and such trigger signal stays high until the PC holds a value matching the end triggerpoint, thus toggling the trigger signal down to 0. Two breakpoints are also leveraged, coupled with the triggerpoints, to manage the oscilloscope’s behavior. A breakpoint is configured to an address that is reached before the start triggerpoint one is employed to interrupt the CPU computation in order to get the oscilloscope ready to perform the acquisition, while another breakpoint is instead assigned to an address being encountered after the acquisition time window to interrupt again the CPU computation and dump the collected data that is stored in the oscilloscope’s buffer.
Such combined usage of the breakpoints and triggerpoints in the prototype execution scenario is accurately replicated in the simulation one for two main purposes. On the one hand, it enables a perfect match from the temporal point of view between the switching activity VCD collected in the simulation and the power trace acquired by the oscilloscope during the prototype execution. On the other hand, it makes it possible to collect the VCD solely for the time window of interest to the SCA analysis, thus limiting the simulation’s execution time. The latter is a particularly critical aspect due to the need to perform a timing simulation of a post-route netlist to have maximum correspondence with the design prototyped on the board and thus collect meaningful switching activity statistics.
Algorithm 1 Chaff SCA countermeasure [40].
1:function ChaffEncrypt (cipher, key, ptx, numChaff)
2:     mainThread $\leftarrow$ CREATE (cipher, key, ptx)
3:     chaffKey $\leftarrow$ GENERATECHAFFKEYS (key, numChaff)
4:     for i in 1 to numChaff do
5:         chaffThreads[i] $\leftarrow$ CREATE (cipher, chaffKey[i], ptx)
6:     STARTRANDOMSCHEDULER (mainThread, chaffThreads)
7:     while ISRUNNING (mainThread) do
8:         WAIT ()
9:     for i in 1 to numChaff do
10:         KILL (chaffThreads[i])
V-C Experimental analysis
The experimental analysis applies to the computing platform configured as described in Section V-B the SCA countermeasure and attack techniques from the state of the art included in the framework, quantitatively assessing their effectiveness.
V-C1 SCA countermeasure techniques
The SCA countermeasures include microarchitecture-level ones, such as clock frequency randomization, and architecture- and OS-level ones, such as the morphing and chaff techniques.
Clock frequency randomization
The DFS actuator can be used to continuously vary, in a random fashion, the frequency of the clock signal fed to the CPU, as previously described in Section IV-C, with the purpose of producing a distortion in the power consumption trace and therefore reducing the information leakage [15]. The DFS, when in random reconfiguration mode, uses indeed as its target frequencies the values generated by the internal TRNG, rather than setting them through the debug infrastructure. The DFS actuator produces a clock signal that can be reconfigured to up to 1024 different frequencies every few tens of microseconds and without any glitching and gating effects, enabling a large variability and ensuring the effectiveness of the countermeasure.
Morphing
The morphing countermeasure leverages the TRNG to randomly modify the execution of a cryptographic operation [31]. In the use case considered in the experimental evaluation, morphing targets the AddRoundKey and SubBytes steps of the AES cryptosystem. Each AddRoundKey execution makes use of a randomly selected version of the XOR operation, chosen among 8 equivalent implementations with different power consumption profiles. SubBytes executions are instead morphed by employing two S-Boxes that are masked by random values and periodically refreshed.
Chaff
The TRNG component and the support for software multithreading provided by FreeRTOS enable the adoption of the chaff approach, described in [40] and whose pseudocode is listed in Algorithm 1. The latter technique executes, concurrently to a thread for the actual instance of a cipher (line 2 in Algorithm 1), e.g., the AES encryption of a plaintext with a certain key, a set of additional threads performing the same cryptographic operation on the same input but with different keys, properly generated to be correlated with the original one (lines 3–5). All the threads are started and executed in parallel, according to a random scheduler that leverages the TRNG instantiated in the SoC (line 6), until the thread performing the actual cryptographic operation has completed (lines 7–10).
V-C2 SCA attack techniques
The SCA attacks provided out of the box by the framework include 1) a CPA attack [7], that extracts the secret key of the target cryptosystem executions by correlating their power signature with their operating behavior, performed in the experimental campaign on up to 1 million AES power traces, 2) a template profiled attack [9], that leverages the Bayes’ theorem to estimate the probability of a key given its multivariate normal distribution, carried out by using 1024 traces for each of the 256 possible AES key bytes during both the profiling and attack phases, and 3) an ML-based attack leveraging convolutional neural networks (CNNs) [32], that learn a recurrent pattern of the leakage to predict the secret key and whose training and inference are carried out on the same dataset as the template attack.
V-C3 SCA security assessment
The experimental analysis assesses the vulnerability of the computing platform to the previously described attacks while executing the AES cryptosystem, whose source code is included in the framework. In particular, we evaluate whether each attack, specifically targeting the SubBytes step of the first AES round, succeeds or fails when executing the plain AES cryptosystem without any protection, a software-masked version of AES, and plain AES protected by the clock frequency randomization, morphing, and chaff SCA hiding countermeasures.
TABLE II: Quality metrics obtained by attacks on the experimental platform when implementing different countermeasures. Legend: – undefined.
Attack
Countermeasure Quality metric CPA Template CNN
None Guessing entropy 1 1 1
Guessing distance 0.86 0.77
Success rate 100% 100% 100%
Number of traces 180 3 10
Software masking Guessing entropy 108.75 153.22 1.06
Guessing distance -0.57 0.45
Success rate 0% 0% 98%
Number of traces
Clock frequency Guessing entropy 112.75 127 72.20
randomization Guessing distance -0.52 -0.31
Success rate 0% 0% 6%
Number of traces
Morphing Guessing entropy 1 1 1
Guessing distance 0.38 0.81
Success rate 100% 100% 100%
Number of traces 3072 580 8
Chaff Guessing entropy 105.25 135 122
Guessing distance -0.52 -0.48
Success rate 0% 0% 0%
Number of traces
The attacks’ effectiveness is evaluated according to 1) the guessing entropy, defined as the average rank position of the correct key among all possible key guesses, 2) the guessing distance, which represents the normalized probability distance between the correct key and the first-ranked non-correct one, 3) the success rate, i.e., the percentage of attacks that succeed in delivering the secret key, and 4) the minimum number of traces required to obtain a prediction that is always correct. TABLE II lists the values of such four quality metrics obtained for the various countermeasure-attack combinations considered in the experimental campaign. In particular, the attacks that always succeed, i.e., with guessing entropy equal to 1 and 100% success rate, respectively, list the corresponding number of traces employed to achieve such a perfect attack outcome. The execution of the plain AES application is successfully attacked by all three considered techniques, while the masked version is broken only by the CNN attack. The clock frequency randomization and chaff countermeasures prove instead to be effective against the three SCA attacks, whereas applying morphing does not protect the computing platform from any of them.
More in detail, clock frequency randomization and chaff are the most effective countermeasures against the proposed SCA attacks, as they achieve a high guessing entropy and a success rate close to 0% in every scenario. The morphing countermeasure is instead shown to be more vulnerable, as every technique can break it. The CNN attack can isolate the correct key with a guessing distance of 0.81, using only eight traces, while CPA and template attacks are less effective though successful. Guessing distance values are notably undefined for the CPA attack, since it is not a profiled attack.
Figure 6 depicts how the SCA attacks against the execution of plain AES evolve as the number of traces increases when applying no SCA countermeasure. The CPA attack, as shown in Figure 6(a), requires 180 traces to correctly identify the correct key. More traces are needed to reduce the correlation coefficient of the wrong key guesses, drawn in red, whereas the Pearson correlation coefficient (PCC) of the correct key, drawn in green, remains steadily around 0.43. Figure 6(b) demonstrates instead how the guessing entropy improves in the template and CNN attacks as more traces get analyzed and how it stabilizes to 1 after 3 and 10 traces have been processed, respectively.
VI Conclusions
This manuscript introduced a novel open-source framework for research on SCA targeting FPGA-based IoT/microcontroller-class computing platforms. The framework includes a RISC-V-based IoT-class SoC that features an ad-hoc debug infrastructure to maximize the observability and controllability of the computing platform and thus simplify the execution of SCA attacks, as well as a DFS actuator, a TRNG, and a timer that provide support for a set of state-of-the-art SCA countermeasures available out of the box. A complete automated flow encompasses the configuration of the SoC, the execution of target applications and corresponding collection of side-channel information, and the analysis to identify eventual SCA vulnerabilities and pinpoint the sources of side-channel information leakage.
(a) Pearson correlation coefficient (PCC) of CPA attack
(b) Guessing entropy of template and CNN attacks
Figure 6: SCA attacks against unprotected AES.
The user is encouraged and empowered to expand the capabilities of the hardware-software infrastructure and support novel SCA attacks and countermeasures by the open-source nature of the framework, its adoption of standard languages for both its hardware and software components, and the usage of widely available devices and tools.
Future developments foresee the addition of a multi-core CPU architecture and cryptographic hardware accelerators and support for fault attacks and countermeasures against them.